# SCMPPI: Supervised Contrastive Multimodal Framework for Predicting Protein-Protein Interactions

### v2

@@ 2 Related Works
In this study, we apply supervised contrastive learning [15]to the PPI task. By pulling positive sample pairs closer, pushing negative sample separation farther apart,and extracting label information, we have improved the ability of the multimodal model to embed sequence and structural information. We also incorporate a negative sample filtering mechanism to reduce the likelihood of false negatives and modify the loss of SuoCons $\mathcal{L}^{Sup}_{out}$ to $\mathcal{L}^{P-Sup}$ for PPI. This section delineates the architecture of the SCMPPI framework, an innovative concept devised to facilitate efficient and high-quality PPI prediction.The structural and design principles of the framework are meticulously expounded in Section 3.1, followed by an overview of the application of multimodal fusion techniques in Section 3.2, and finally its core contrast learning module (Section 3.3) and classifier (Section 3.4), respectively.
@@ 2.6.5 Testing on PPI Network Datasets
In this study, we applied the SCMPPI model to predict PPIs within three key biological networks: CD9, Wnt-related pathway, and Cancer-specific networks. The model successfully identified critical PPIs, revealing biologically and clinically relevant proteins [44, 45].
CD9, a widely expressed tetraspanin protein, was identified as a central node. Its role in tumor progression is context-dependent—exhibiting tumor-suppressive effects in cancers like prostate cancer, but linked to poor prognosis in others such as pancreatic cancer [46].
In the Wnt-related pathway, SCMPPI highlighted AXIN1 and WNT9A, both central to Wnt signaling regulation. AXIN1 acts to suppress the pathway by promoting $\beta$-catenin degradation, while WNT9A activates it as a ligand. Dysregulation of this pathway is strongly linked to cancer progression [47].
For the Cancer-specific network, CDK1 and TP53 were predicted as key proteins. CDK1 facilitates cell cycle progression, while TP53, a well-known tumor suppressor, controls DNA repair and apoptosis. Mutations in TP53 are commonly implicated in cancer development [48, 49].
Overall, SCMPPI effectively captures meaningful biological insights and provides valuable support for cancer-related research and therapeutic development.

### v3

@@ 2 Related Works
In this study, we apply supervised contrastive learning [15]to the PPI task. By pulling positive sample pairs closer, pushing negative sample separation farther apart,and extracting label information, we have improved the ability of the multimodal model to embed sequence and structural information. We also incorporate a negative sample filtering mechanism to reduce the likelihood of false negatives and modify the loss of SuoCons $\mathcal{L}^{Sup}_{out}$ to $\mathcal{L}^{P-Sup}$ for PPI.
3 Approach
This section delineates the architecture of the SCMPPI framework, an innovative concept devised to facilitate efficient and high-quality PPI prediction.The structural and design principles of the framework are meticulously expounded in Section 3.1, followed by an overview of the application of multimodal fusion techniques in Section 3.2, and finally its core contrast learning module (Section 3.3) and classifier (Section 3.4), respectively.
@@ 3.4 Prediction Module
4 Experiments
In this section, we describe the experimental setup we used and analyse what we found.
@@ 4.2.5 Testing on PPI Network Datasets
In this study, we applied the SCMPPI model to predict PPIs within three key biological networks : CD9, Wnt-related pathway, and Cancer-specific networks. The model successfully identified critical PPIs in the Figure 9, revealing biologically and clinically relevant proteins [44, 45].
Firstly, CD9 is a tetraspanin transmembrane protein widely distributed across various cell types, involved in processes such as cell adhesion, migration, and fusion source[46]. In tumor metastasis, CD9 plays a dual role. In some cancers, such as prostate cancer, high expression of CD9 is associated with lower metastatic ability, exhibiting a tumor-suppressive effect. However, in other types of cancer, such as pancreatic cancer, high CD9 expression correlates with poor prognosis. Therefore, the role of CD9 in tumor prognosis and treatment should be analyzed based on specific cancer types.
The Wnt-related pathway plays a crucial role in embryonic development, cell proliferation, differentiation, migration, and tissue homeostasis. Abnormal activation of this pathway is closely associated with the onset and progression of various cancers source[47]. AXIN1 and WNT9A are core proteins in this pathway. AXIN1 inhibits the activity of the Wnt pathway by forming a complex that degrades $\beta$-catenin, while WNT9A, as a ligand protein, activates the pathway. Regulating the activity of the Wnt signaling pathway holds potential clinical value for cancer prevention and treatment.
Finally, in the Cancer-specific network, CDK1 and TP53 are two key proteins. CDK1 regulates cell proliferation during the G2/M phase of the cell cycle, and its abnormal activation can promote tumor cell proliferation source[48]. TP53, as a tumor suppressor gene, is responsible for DNA repair and apoptosis, and mutations in TP53 are commonly found in various cancers source[49]. Studying these proteins and their interactions can provide an important theoretical foundation for cancer therapy.
Through the accurate predictions of the SCMPPI model, we have not only validated the important roles of these core proteins in various biological processes but also opened up new research directions for the early diagnosis and targeted treatment of cancer.